The score when two fair six-sided dice are thrown is the sum of the two numbers on the upper faces.
- (a)Show that the probability that the score is 4 is 1/12.[1]
- (b)The two dice are thrown repeatedly until a score of 4 is obtained. The number of throws taken is denoted by the random variable X. Find the mean of X.[1]
- (c)Find the probability that a score of 4 is first obtained on the 6th throw.[1]
- (d)Find P(X < 8).[2]
Mark scheme
Marking points from the official mark scheme (May/June 2020). The mark or mark type is shown in brackets.
- [1] Prob of 4 (from 1,3, 3,1 or 2,2) = 3/36 = 1/12 AG
- [1] Mean = 1/(1/12) = 12
- [1] (11/12)^5 x 1/12 = 0.0539 or 161051/2985984
- [1] 1 - (11/12)^7
- [1] 0.456 or 16344637/35831808
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →